OpenStudy (anonymous):

Which group of points lies on the line y = 4x - 3?

OpenStudy (anonymous):

Are there choices to choose from?

OpenStudy (anonymous):

Regardless, you can test points (x1, y1) by substituting in x1 for x in your equation. If you get y1, then that point is on the line. If it is not, then the point is not on the line.

OpenStudy (anonymous):

A (1, 1); (2, –1); (–1, –7), B(1, 1); (2, 5); (–1, –7), C(1, 1); (2, 5); (-1, –1) D(1, 1); (2, –1); (–1, –1)

OpenStudy (anonymous):

I would start by checking (2, -1) to see if that is on the line. y = 4(2) - 3 y = 8 - 3 = 5 Therefore, (2, -1) is NOT on the line.

OpenStudy (anonymous):

Just go through each and check the points.

OpenStudy (anonymous):

But we do know that (2, 5) is on the line.